What volume of 15.9 M Nitric acid would be required to make 3.2 L of 2.00 M nitric acid?
1
Expert's answer
2020-12-18T08:38:44-0500

using M1V1 = M2V2

15.9 × V1 = 3.2×2

V1 = 6.4/15.9 = 0.402 L